OpenStudy (anonymous):

Lily buys a $425 TV and pays $31.88 in tax. If she also buys a $200 DVD recorder, how much additional tax will she pay?

OpenStudy (bahrom7893):

Okay first let's find out how much tax she pays: $425->100% $31.88->x%

OpenStudy (bahrom7893):

x = 31.88*100/425=3188/425=7.5

OpenStudy (bahrom7893):

So she pays 7.5% tax. Now do the following: 200 - 100% x - 7.5%

OpenStudy (bahrom7893):

x = 200*7.5/100 = 2*7.5 = 15

OpenStudy (bahrom7893):

She has to pay $15 more

OpenStudy (anonymous):

wow you did that fast!

OpenStudy (anonymous):

thanks!

OpenStudy (bahrom7893):

I'm a superhero after all haha =)

OpenStudy (anonymous):

i have another one I cant figure out

OpenStudy (bahrom7893):

Post it as a new thread, this gets slow if it gets long

OpenStudy (anonymous):

The local community theater has fixed expenses of $1500 per week. The theater sells about 150 tickets to a play each week when it charges $20 per ticket. When it charges $14 per ticket, about 210 tickets are sold. Write the equation that relates ticket price and number of tickets sold, and use it to predict how many tickets they should expect to sell if they change the ticket price to $9.50.

OpenStudy (anonymous):

thats the one I have posted

OpenStudy (bahrom7893):

hmmmm will have to think about this

OpenStudy (anonymous):

you want to go to the question so I can give you credit?

OpenStudy (bahrom7893):

no ill just do it here

OpenStudy (anonymous):

cool

OpenStudy (bahrom7893):

dang this is a tough one..

OpenStudy (anonymous):

i thought so too

OpenStudy (bahrom7893):

20*150=3000 hmm

OpenStudy (bahrom7893):

14*210=2940.. these are just random thoughts

OpenStudy (bahrom7893):

dang lol this is making me feel bad..

OpenStudy (anonymous):

why?

OpenStudy (bahrom7893):

because I am having trouble solving it..

OpenStudy (bahrom7893):

do u have the answer to it btw?

OpenStudy (anonymous):

i have been on this problem for like 3 hours

OpenStudy (anonymous):

no its for my class

OpenStudy (anonymous):

i was thinking 9.5 * 270=2565

OpenStudy (bahrom7893):

I was thinking: 20 - 150 9.5 - x x = 9.5*150/20 = 71.25

OpenStudy (anonymous):

looks good to me

OpenStudy (bahrom7893):

but then: 14-210 9.5-x x=142.5

OpenStudy (bahrom7893):

that doesn't make any sense

OpenStudy (bahrom7893):

arghhhh

OpenStudy (bahrom7893):

I guess u will have to hang around for our brilliant minds to show up. Those include: joemath(bunch of numbers at the end) amistre satellite alchemista myininaya and imranmeah91

OpenStudy (anonymous):

this is my last problem...i am stummped

OpenStudy (bahrom7893):

Sorry I have to write my own speech right now (for a class), but if this is due on monday or something, i can take a look at the problem later tonight and see if i can solve it.

OpenStudy (anonymous):

so the equation would look like this? 9.50*x=t

OpenStudy (anonymous):

this is due today

OpenStudy (bahrom7893):

actually give me like 15 minutes..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what is your speech on?

OpenStudy (bahrom7893):

Bernard Baruch, anyway working on this problem. This looks simple, like logic, there's actually like a lot more info given than necessary which makes it harder..

OpenStudy (anonymous):

i agree

OpenStudy (bahrom7893):

can u check the answer if i give it to you? like is this an internet assignment that will say wrong if u enter the wrong answer?

OpenStudy (anonymous):

yes

OpenStudy (bahrom7893):

YAY! I think i got it.

OpenStudy (anonymous):

actuall I am taking collage math on line

OpenStudy (bahrom7893):

how many trials do u get?

OpenStudy (anonymous):

not by choice

OpenStudy (anonymous):

but i have a site where I can check answers

OpenStudy (anonymous):

i just need to find the equation

OpenStudy (bahrom7893):

yea im tryin..

OpenStudy (anonymous):

i came up with this but not sure how to apply it Simplifying x * 9.5 = b Reorder the terms for easier multiplication: 9.5x = b Solving 9.5x = b Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Divide each side by '9.5'. x = 0.1052631579b

OpenStudy (bahrom7893):

this 2nd one is throwing me off.. their revenue is not constant, it's either 3000 or 2940

OpenStudy (anonymous):

my brain is mush... I will just do 9.5*270=2565

OpenStudy (bahrom7893):

Wait isn't this like a line?

OpenStudy (bahrom7893):

f(20)=150 f(14)=210

OpenStudy (anonymous):

now you lost me

OpenStudy (anonymous):

It is a line. The 'fixed expenses' part doesn't seem relevant. You are given two points and asked to find the line that relates them

OpenStudy (anonymous):

how do we do that

OpenStudy (anonymous):

Yes exactly Bahrom.

OpenStudy (bahrom7893):

oh okay working on it

OpenStudy (bahrom7893):

Let: y=150, x = 20 y1 = 210, x1 = 14

OpenStudy (anonymous):

You have two points (14,210) and (20,150). Plug them into the slope formula to find the slope of the line. Then use the slope and one of your points in the point-slope formula for a line.

OpenStudy (bahrom7893):

Slope of that line is: M = (y-y1)/(x-x1) = (150-210)/(20-14) = -60/6=-10

OpenStudy (anonymous):

Once you have the equation for the line, you'll want to plug in 9.5 for x and see what your y value is to find the number of tickets sold.

OpenStudy (bahrom7893):

y-y1 = M(x-x1) y-210=-10(x-14) y = -10x + 140 + 210 y = -10x +350

OpenStudy (bahrom7893):

That's the equation where y is the # of tickets and x is the price.

OpenStudy (bahrom7893):

So when price is 9.5 y = -10*9.5+350 = -95+350=350-95=255

OpenStudy (bahrom7893):

So the amount of tickets they would sell if the price was 9.5 would be 255 and the equation would be: y = -10x+350, where y is the # of tickets and x is the price.

OpenStudy (bahrom7893):

Check the answer, and thanks Polpak!

OpenStudy (anonymous):

great job all thanks for all your help bahrom! good luck on your speech!

OpenStudy (bahrom7893):

lol np =)

OpenStudy (anonymous):

it makes sence to me now!